Jetson Spacecraft Corp. shows the following information on its 2009 income statement sales $213,000; costs $89,000; other expenses $6,000; depreciation expense- $9,100; interest expense = $13,500; taxes = $28,620; dividends = $9,400. In addition, youre told that the firm issued $7,500 in new equity during 2009 and redeemed $9,100 in outstanding long-term debt. (a)What is the 2009 operating cash flow? (Click to select) (b)What is the 2009 cash flow to creditors? (Click to select) v (c)What is the 2009 cash flow to stockholders? (Click to select) ▼ If net fixed assets increased by $18,000 during the year, what was the addition to NWC? Click to select)
Answer #1

1: Operating cash flow

Sales 213000

Less: costs 89000

Other expenses 6000

Taxes 28620

Operating cash flow = 89380

2: Cash flow to creditors = Interest + Debt redeemed

= 13500+9100

= 22,600

3: Cash flow to stockholders = Dividend – Equity issued

= 9400-7500

= 1900

4: Cash flow from assets = Cash flow to creditors + cash to stockholders

= 22600+1900

= 24,500

=OCf – change in NWC _ net capital exp

24500 = 89380- Change in NWC – 18000

Change in NWC= 89380-18000-24500

=46880